Feed the Hungry Foundation

Feed the Hungry Foundation is a non-profit organization that has a cost of capital of 13 percent. The foundation is considering the replacement of a piece of equipment. The old machine has a book value of $6,000 and a remaining estimated life of 5 years with no salvage value at that time. The value of the old machine is currently $3,000. The new equipment will cost $12,000. It has an estimated life of 5 years with no salvage value. Annual cash operating costs are $4,500 for the old machine and $2,100 for the new machine (ignore tax impact).

Note: ignore any tax impacts

  4.1.   Refer to Feed the Hungry Foundation. What is the present value of the operating cash outflows for the old machine?

  4.2.   Refer to Feed the Hungry Foundation. What is the present value of the operating cash outflows for the new machine?

  4.3.   Refer to Feed the Hungry Foundation. What is the present value of salvage value of the old machine if it is replaced now?

  4.4.   Refer to Feed the Hungry Foundation. Would you advise the organization to replace the machine?
Answer #1

4.1 Present value of the operating cash outflows for the old machine = $4500 * 3.5172 (PVF 13%, 5) = $15827

4.2 Present value of the operating cash outflows for the new machine = $2100 * 3.5172 (PVF 13%, 5) = $7386

4.3 Present value of salvage value of the old machine if it is replaced now = $3000

4.4 Net Present value of machine = Net cash outflow - Present value of saving

= (12000-3000) - (15827 - 7386)

= - $559

The organization should not replace the machine because it has negative NPV.